Mrs. Anderson had twice as many chickens as ducks

How many chickens did she have in the beginnning?

Mrs. Anderson had twice as many chickens as ducks. She sold 272 chickens and 16 ducks. She then had half as many chickens as ducks. How many chickens did she have in the beginning?
Answer #1

Let c = chickens and d = ducks, then:

c = 2d

c - 272 = 1/2 (d - 16)

We can rewrite the expression in terms of one variable, d:

2d - 272 = 1/2 (d - 16)

2d - 272 = 1/2d - 8

3/2d = 264

d = 176

c = 2 * 176 = 352

answered by: mathpower